Program Information
Recent Log Entries
- FML SAM Update
- ACG Report 01/26/2026
- ATLis Comments Daily Digest
- CIS Work Summary 01/23/2026
- iocse20 restored
- iochlc2 restored
- SOUTH_LINAC has been disabled in alarm handler
- SOUTH_LINAC has been enabled in alarm handler
- Follow-up Re: Hall D Dumplet In/Out State Indeterminant
- Program Deputy Summary, presented 2026-01-26
- MOLLER Install Update 01/23/2026
- ODH head OM203 at LERF Tunnel - Zone 02 has been bypassed.
- ODH head OMB06 at Hall B - Access Tunnel has been bypassed.
- ATLis Comments Daily Digest
- ODH head OM06 (near access door 2) in Hall B appears to be failing











